# Anomalous Ionization in the Central Molecular Zone by sub-GeV Dark Matter

### v1

@@ Abstract
We demonstrate that the anomalous ionization rate observed in the Central Molecular Zone can be attributed to MeV dark matter annihilations into $e^{+}e^{-}$ pairs for galactic dark matter profiles with slopes $\gamma>1$. The low annihilation cross-sections required avoid cosmological constraints and imply no detectable inverse Compton, bremsstrahlung or synchrotron emissions in radio, X and gamma rays. The possible connection to the source of the unexplained $511$ keV line emission in the Galactic Center suggests that both observations could be correlated and have a common origin.
I introduction
Various measurements are known to indicate an unexpectedly high rate of H2 ionization rates around the Galactic Center (GC), particularly in the Central Molecular Zone (CMZ). This has been evidenced by a variety of tracers, including methanol production Coutens et al. (2017), (indicating an ionization rate of $\zeta\sim 10^{-15}$ s-1), and the observation of $\textrm{H}_{3}^{+}$ lines Oka and Geballe (2020); Indriolo et al. (2009); Goto et al. (2014) (requiring an ionization rate around $\zeta\sim 2\times 10^{-14}$ s-1Le Petit et al. (2016); Ravikularaman et al. (2024); Geballe and Oka (2010) and other ionized molecules, such as $\textrm{H}_{3}$+ Indriolo et al. (2015); van der Tak et al. (2006) (indicative of an ionization rate $\gtrsim 10^{-15}$ s-1).
Recently, Ref. Ravikularaman et al. (2024) showed that the diffuse flux of cosmic rays (CRs) falls short of explaining such high ionization rates in the CMZ, arguing that it requires an extremely large injection power for CRs at the GC, of around $10^{40}-10^{41}$ erg s-1 (see also Ref. Phan et al. (2018)). Moreover, it was shown that such a CR flux would exceed the MeV gamma-ray emission observed in the inner galaxy ($|l|<$ $30^{\circ}$ and $|b|<$ $15^{\circ}$). This may indicate the presence of a source highly concentrated around the GC that is producing low energy CRs that lose their energy very quickly, via ionization of molecular gas Indriolo et al. (2009). From the latter, one can infer that these particles must be injected at low energies, to efficiently be confined around the CMZ thereby preventing them from propagating long distances.
@@
Figure 1: Maximum ionization rate that can be provided by sub-GeV DM, allowed by CMB constraints on $\langle\sigma v\rangle$ Lopez-Honorez et al. (2013); Slatyer (2016). We compare the predicted ionization rate obtained with an analytical calculation and with a numerical calculation using a dedicated CR propagation code (DRAGON2). The shaded area represents the ionization rate observed in the CMZ.
@@
The motivation for sub-GeV DM is strong, and has previously been discussed in detail in the literature, see e.g. Refs. Battaglieri et al. (2017); Cirelli et al. (2021); Balan et al. (2024a) and Ref. Cirelli et al. (2024) for a recent review. Presently, CMB constraints Lopez-Honorez et al. (2013); Slatyer (2016), provide the most robust constraints for annihilation of sub-GeV DM particles, and, as shown in Fig. 1, these are consistent with a DM origin of the anomalous ionization rate observed in the CMZ for profiles generally cuspier than a NFW DM profile Navarro et al. (1996) and low DM masses (and even allow for combinations of mass and DM profiles that could exceed the observed ionization rate), due to the fact that the (low-energy) DM products can deposit roughly all their energy within the CMZ. Here we investigate whether Galactic probes are able to constrain and rule out this explanation of the CMZ ionization rate.
@@
Among the most interesting probes for sub-GeV DM is the $511$ keV line emission, that forms from the decay of positronium formed by the positrons in the Galaxy, and was shown to produce strong constraints on sub-GeV DM annihilation into Standard Model states De la Torre Luque et al. (2023a); Vincent et al. (2012). In fact, $511$ keV line emission from the bulge of the Galaxy remains a mystery Prantzos et al. (2011); Kierans et al. (2019); Siegert (2023) and sub-GeV DM was proposed as a possible explanation Boehm et al. (2004). The observation of the line at the bulge requires a steady production of positrons at a rate of a few times $10^{50}$ $e^{+}$/yr Siegert et al. (2016); Kierans et al. (2020). Once the electrons and positrons are produced, they lose energy, mainly via ionization of the molecular gas in the CMZ. If the medium is dense enough, as is the case for the CMZ, they release all their energy into the medium after travelling less than $\sim 100$ pc (for $e^{\pm}$ with energies below a few tens of MeV for standard diffusion parameters found in CR analyses). Once thermalized, the relativistic positrons produce positronium bound states, in the form of a singlet state, called para-positronium (p-ps), that decays into two gamma rays each with $511$ keV energy, or in a triplet state, called ortho-positronium (o-ps), that decays in $3$ photons Badertscher et al. (2007) producing continuum emission below $511$ keV. We will also discuss the possibility that both observations are connected and explore whether the parameter space of sub-GeV DM that can explain both effects is currently ruled out.
There is another long-standing anomalous emission from the GC that has been related to DM Goodenough and Hooper (2009); Hooper and Goodenough (2011) and other leptonic sources – from milisecond pulsars (MSPs) Abazajian and Kaplinghat (2012); Yuan and Zhang (2014) to leptonic outbursts Cholis et al. (2015): The GeV Fermi-LAT Galactic Center Excess (GCE) Ackermann et al. (2017). This excess spatially coexists with the CMZ (although it is more extended) and its distribution can be well fitted with a contracted NFW DM profile Di Mauro (2021) (c-NFW). The analyses of Fermi-LAT data suggest a luminosity of $10^{37}$ erg s-1 to explain the GCE. As we discuss in this work, it is quite unlikely that the source of this excess can be linked to the other anomalies described above, but understanding these anomalous observations can allow us to understand better the dynamics of the GC.
@@
The remainder of this paper is as follows: We first detail our estimates of the CMZ ionization rate from the annihilation of sub-GeV DM particles. Then, we test if the observations of the CMZ ionization rate are compatible with cosmological constraints and if there would be any associated significant inverse Compton or bremsstrahlung emission. Finally, we discuss whether the CMZ ionization rate and the $511$ keV line emission, could share a common (DM annihilation) origin and comment on the potential of finding correlations between these two observables.
II Ionization rate from sub-GeV DM annihilation in the CMZ
@@
The CMZ, located around the GC, is an approximately cylindrical region around $200$ pc wide and $100$ pc high Ferriere et al. (2007); Ravikularaman et al. (2024), of high-density molecular gas ($n_{\textrm{H}}\sim 150$ cm-3). Low energy CRs $\lesssim 100$ MeV are subject to severe ionization energy losses that limit their propagation (see Fig. 3). At higher energies, these particles can escape the CMZ and produce potentially detectable secondary diffuse emission (such as bremsstrahlung, inverse Compton or synchrotron radiation).
@@
where we the factor $2$ accounts for both $e^{-}$ and $e^{+}$, $E_{\textrm{max}}$ is the maximum kinetic energy with which the electrons or positrons are injected and $E_{\textrm{min}}$ is the minimum kinetic energy to produce ionization in the gas ($13.6$ eV), $\sigma(E)$ is the $e$-H2 cross-section, calculated as in Eq. (7) of Ref. Padovani et al. (2009), $\theta_{e}(E)$ is the number of secondary ionizations per primary ionization, calculated as Eq. (2.23) of Ref. Krause et al. (2015), and $J(E)$ is the flux that can be computed as
@@
where $\beta(E)c$ is the speed of the electron or positron and $\frac{d\phi_{e}}{dE}$ is their differential flux in the CMZ. We compute the differential flux of electrons and positrons induced by sub-GeV DM annihilation in two ways: an analytical approximation that allows us to estimate the diffuse $e^{\pm}$ flux within the CMZ, and a numerical calculation that allows us to compute their all-sky diffuse emission. We also note that these annihilations are expected to also produce ionizing photons through final state radiation, however, the cross-section for ionizing H2 is $\sim 10^{-6}$ times lower than from electrons/positrons at $1$ MeV, and negligible in comparison with the ionization caused by $e^{\pm}$.

### v3

@@ Abstract
We demonstrate that the anomalous ionization rate observed in the Central Molecular Zone can be attributed to MeV dark matter annihilations into $e^{+}e^{-}$ pairs for galactic dark matter profiles with slopes $\gamma>1$. The low annihilation cross-sections required avoid cosmological constraints and imply no detectable inverse Compton, bremsstrahlung or synchrotron emission in radio, X and $\gamma$ rays. The possible connection to the source of the unexplained $511$ keV line emission in the Galactic Center suggests that both observations could be correlated and have a common origin.
preprint: KCL-2024-52
Figure 1: Top panel: Maximum ionization rate that can be provided by sub-GeV DM (i.e. the average ionization rate within the CMZ obtained from the upper limits on $\langle\sigma v\rangle$ constraints allowed by CMB Lopez-Honorez et al. (2013); Slatyer (2016)). We compare the predicted ionization rate obtained with an analytical calculation and with a numerical calculation using a dedicated CR propagation code (DRAGON2Evoli et al. (2017, 2018); De La Torre Luque (2023). The shaded area represents the ionization rate observed in the CMZ Ravikularaman et al. (2024); Miller et al. (2020); Indriolo et al. (2015); Rivilla et al. (2022); Yusef-Zadeh et al. (2012, 2007); Oka et al. (2005); Goto et al. (2011); Geballe and Oka (2010); Le Petit et al. (2016); Sanz-Novo et al. (2024). Bottom panel: Radial profile of the ionization rate expected from DM annihilation ($\langle\sigma v\rangle=10^{-35}$ cm3/s) for the Moore (solid line, multiplied by $10^{-7}$ to have an easy comparison) and NFW (dashed lines) profiles. The different colors indicate different DM masses.
Introduction— Various measurements are known to indicate an unexpectedly high rate of H2 ionization rates around the Galactic Center (GC), particularly in the Central Molecular Zone (CMZ). This has been evidenced by a variety of tracers, including methanol production Coutens et al. (2017), (indicating an ionization rate of $\zeta\sim 10^{-15}$ s-1), and the observation of $\textrm{H}_{3}^{+}$ lines Oka and Geballe (2020); Indriolo et al. (2009); Goto et al. (2014) (requiring an ionization rate around $\zeta\sim 2\times 10^{-14}$ s-1Le Petit et al. (2016); Ravikularaman et al. (2024); Geballe and Oka (2010) and other ionized molecules, such as $\textrm{H}_{3}$+ Indriolo et al. (2015); van der Tak et al. (2006) (indicative of an ionization rate $\gtrsim 10^{-15}$ s-1).
Recently, Ref. Ravikularaman et al. (2024) showed that the diffuse flux of cosmic rays (CRs) falls short of explaining such high ionization rates in the CMZ, arguing that it requires an extremely large injection power for CRs at the GC, of around $10^{40}$-$10^{41}$ erg s-1 (see also Ref. Phan et al. (2018)). This may indicate the presence of a source highly concentrated around the GC that is producing low energy CRs that lose their energy very quickly, via ionization of molecular gas Indriolo et al. (2009). From the latter, one can infer that these particles must be injected at low energies, to be efficiently confined around the CMZ thereby preventing them from propagating long distances.
@@
The motivation for sub-GeV DM is strong and has previously been discussed in detail in the literature, see e.g. Refs. Battaglieri et al. (2017); Cirelli et al. (2021); Balan et al. (2024a) and Ref. Cirelli et al. (2024) for a recent review. Presently, CMB constraints Lopez-Honorez et al. (2013); Slatyer (2016), provide the most robust constraints for annihilation of sub-GeV DM particles, and, as shown in Fig. 1, these are consistent with a DM origin for the anomalous ionization rate observed in the CMZ for profiles generally cuspier than a NFW DM profile Navarro et al. (1996) and low DM masses (and even allow for combinations of mass and DM profiles that could exceed the observed ionization rate), due to the fact that the (low-energy) DM products can deposit roughly all their energy within the CMZ. Here we investigate whether Galactic probes are able to constrain and rule out this explanation of the CMZ ionization rate .
@@
The $511$ keV line emission from the bulge of the Galaxy remains a mystery Prantzos et al. (2011); Kierans et al. (2019); Siegert (2023) and sub-GeV DM was proposed as a possible explanation Boehm et al. (2004). The emission can be used to strongly constrain sub-GeV DM, which injects Standard Model particles that yield electrons and positrons, forming positronium that can decay in the Galaxy De la Torre Luque et al. (2023a); Vincent et al. (2012). The observation of the line at the bulge requires a steady production of positrons at a rate of a few times $10^{50}$ $e^{+}$/yr Siegert et al. (2016); Kierans et al. (2020). Once the electrons and positrons are produced, they lose energy, mainly via ionization of the molecular gas in the CMZ. If the medium is dense enough, as is the case for the CMZ, they release all their energy into the medium after travelling less than $\sim 100$ pc (for $e^{\pm}$ with energies below a few tens of MeV for standard diffusion parameters found in CR analyses). Once thermalized, the relativistic positrons produce positronium bound states, in the form of a singlet state, called para-positronium (p-ps), that decays into two $\gamma$ rays each with $511$ keV energy, or in a triplet state, called ortho-positronium (o-ps), that decays into $3$ photons Badertscher et al. (2007) producing continuum emission below $511$ keV. We will also discuss the possibility that both observations are connected and explore whether the parameter space of sub-GeV DM that can explain both effects is currently ruled out.
@@
The remainder of this paper is as follows: we first detail our estimates of the CMZ ionization rate from the annihilation of sub-GeV DM particles. Then, we test whether the observations of the CMZ ionization rate are compatible with cosmological constraints and if there would be any associated significant inverse Compton or bremsstrahlung emission. Finally, we discuss whether the CMZ ionization rate and the $511$ keV line emission, could share a common (DM annihilation) origin and comment on the potential of finding correlations between these two observables.
@@
Ionization rate from sub-GeV DM annihilation in the CMZ – The CMZ, located around the GC, is an approximately cylindrical region around $200$ pc wide and $100$ pc high Ferriere et al. (2007); Ravikularaman et al. (2024), of high-density molecular gas ($n_{\textrm{H}}\sim 150$ cm-3). Low energy CRs $\lesssim 100$ MeV are subject to severe ionization energy losses that limit their propagation (see Fig. S1 in Supplementary material). At higher energies, these particles can escape the CMZ and produce potentially detectable secondary diffuse emission (such as bremsstrahlung, inverse Compton or synchrotron radiation).
@@
where the factor $2$ accounts for ionization by both $e^{-}$ and $e^{+}$, $E_{\textrm{max}}$ is the maximum kinetic energy with which the electrons or positrons are injected and $E_{\textrm{min}}$ is the minimum kinetic energy to produce ionization in the gas ($15.43$ eV Padovani et al. (2009); Phan et al. (2023)), $\sigma(E)$ is the $e$-H2 cross-section, calculated as in Eq. (7) of Ref. Padovani et al. (2009), $\theta_{e}(E)$ is the number of secondary ionizations per primary ionization, calculated as Eq. (2.23) of Ref. Krause et al. (2015), and $J(E,\textbf{x})$ is the flux that can be computed as
@@
where $\beta(E)c$ is the speed of the electron or positron and $\frac{d\phi_{e}}{dE}$ is their differential flux in the CMZ. We compute the differential flux of electrons and positrons induced by sub-GeV DM annihilation in two ways: (i) an analytical approximation that allows us to estimate the diffuse $e^{\pm}$ flux within the CMZ, and (ii) a numerical calculation that allows us to compute their all-sky diffuse emission. We notice that both approaches used to compute the ionization rate are approximate, given the difficulties of treating diffusion in the CMZ, however, as we show below, the parameter space accessible to explain the measured ionization rate is so large that our conclusion should remain robust. We refer the reader to Section S.I of the Supp. Material for details on these calculations, which includes Refs. De la Torre Luque et al. (2023b, c, d, a); Ginzburg and Syrovatskii (1969); Crank and Nicolson (1947); Akiyama et al. (2022); Aguilar et al. (2021, 2018, 2013); Webber et al. (2002); Stone et al. (2013); Evoli et al. (2020); Cirelli et al. (2011). We also note that these annihilations are expected to also produce ionizing photons through final state radiation, however, the cross-section for ionizing H2 is $\sim 10^{-6}$ times lower than from $e^{\pm}$ at $1$ MeV.
Results — The main constraints on sub-GeV DM arise from cosmological observations of the early Universe such as the effective number of relativistic species $N_{\textrm{eff}}$ and the requirement for successful Big Bang Nucleosynthesis (BBN). These observations can robustly rule out DM masses below around $1$ MeV Boehm et al. (2012, 2013); Nollett and Steigman (2015), and even slightly above, depending on certain assumptions. However, recent work, such as Refs. Escudero (2019); Sabti et al. (2020), has shown compatibility with observation for DM masses above a few MeV, as long as the dominant early Universe channel for DM annihilation is into neutrinos. For higher masses, there are strong astrophysical and cosmological constraints (see a recent review Ref. Balan et al. (2024b) or a more comprehensive review of DM candidates Ref. Cirelli et al. (2024)).
[…]
In terms of constraints from Galactic probes, we remark that X-ray constraints are competitive with those from CMB, although for DM particles with masses below $\sim 10$-$20$ MeV, diffuse reacceleration of the $e^{\pm}$ pairs injected by DM De la Torre Luque et al. (2023d); Seo and Ptuskin (1994) induces uncertainties at MeV energies. We have tested the X-ray emission produced by the DM-induced $e^{\pm}$ population using the Hermes code Dundovic et al. (2021) (as in Ref. De la Torre Luque et al. (2023d)), with the same propagation parameters as used for the estimations of the ionization rate in the CMZ, finding that XMM-Newton observations cannot rule out DM with masses below a few tens of MeV, as shown in the right panel of Fig. S4 (Supp. Material, which contains the additional Refs. Foster et al. (2021); Cirelli et al. (2023b)). We have also calculated the bremsstrahlung emission (using the Hermes code as in Ref. Luque et al. (2024)) at the MeV scale. Although bremsstrahlung emission is expected to be more important for low DM masses Cirelli et al. (2013) and dominates over the inverse Compton emission for DM masses below a few tens of MeV Bartels et al. (2017), we find that this emission is not expected to be significant for the annihilation strength needed to explain the CMZ ionization rate and it does not represent an important constraint (see right panel of Fig. S4 in the SM, which contains the additional Ref. Bouchet et al. (2010)). As expected, the bremsstrahlung emission is not competitive with observables such as the $511$ keV line, X-rays or the CMB. In this context, there are indications that point to an excess of diffuse photons at MeV energies Karwin et al. (2023). Annihilation of MeV DM could contribute to the diffuse emission around the GC and help alleviate the current tension with the expected background emission.
Possible connection with the 511 keV line emission – We now investigate the tantalizing possibility that the $511$ keV line measurements around the GC are correlated with the hydrogen ionization rate in the CMZ. Our discussion is motivated by the following:
ii) Ref. De la Torre Luque et al. (2023a) showed that, due to the propagation of positrons, the longitude profile of the $511$ keV line emission can be reproduced only for DM profiles cuspier than the NFW, also favoring low DM masses (see the Appendix of Ref. De la Torre Luque et al. (2023a) and further discussions in Ref. De la Torre Luque et al. (2024a)). Although a contracted NFW distribution would better fit the signal, the exact contraction index (or value of the slope, $\gamma$) depends on the mass of the DM particle annihilating (given that propagation induces a mass-dependent behavior of the 511 keV emission profile from DM) and the exact propagation parameters used.
iii) The power density from DM annihilations in the CMZ is compatible with observations of the line emission at the GC for a cross-section $\langle\sigma v\rangle\lesssim 10^{-32}$ cm3/s. For example, a $1$ MeV particle in a Moore DM profile can produce $\sim 10^{52}\leavevmode\nobreak\ \rm erg/s\sim 5\times 10^{50}$$e^{+}/$yr for $\langle\sigma v\rangle\lesssim 10^{-33}$ cm3/s.
To test whether both can be connected, we compute the $511$ keV emission in the bulge (integrated in the inner $8^{\circ}$) from the annihilation of sub-GeV DM for the different DM profiles considered in this work, following the same approach as in Ref. De la Torre Luque et al. (2024b), which is consistent with previous evaluations Calore et al. (2021, 2022); De la Torre Luque et al. (2023b, c); Carenza et al. (2024). Adopting an annihilation cross-section of $\langle\sigma v\rangle=10^{-35}$ cm3/s, we depict the predicted $511$ keV emission in Fig. 2 (left axis – solid lines), where we include a shaded gray band that accounts for the observed emission at the bulge Siegert et al. (2016), to which we add a conservative $50\%$ uncertainty (to account for underestimated systematic uncertainties). From a comparison to the ionization rate at the CMZ (right axis of the figure – dashed lines), we observe that the annihilation cross-section needed to explain the $511$ line emission at the bulge is lower than the annihilation cross-section that explains the CMZ ionization rate, by at least a factor of a few, meaning that a simultaneous explanation of both phenomena seems difficult under the DM hypothesis. This may indicate that the source of the CMZ ionization rate is not symmetrically injecting electrons and positrons, and could instead be dominated by ionizing protons (or only electrons).
However, we note here that while uncertainties from propagation parameters similarly affect both predictions, the estimated $511$ keV emission is also affected by the conditions for positronium production. The formation of positronium from thermal positrons is highly dependent on the composition, temperature and density of the medium Guessoum et al. (1991), hindering our ability to make a robust prediction of this quantity (especially around the GC). Therefore, it would be difficult to firmly rule out the DM hypothesis of the CMZ ionization rate from the estimated $511$ keV line bulge emission given that uncertainties in this estimate are larger than an order of magnitude and that the cross-sections needed to match both observations can be close (e.g. for the cNFW profile case with 1 MeV DM where the 511 keV emission is compliant at $\phi^{511\,\textrm{keV}}_{8^{\circ}}\gtrsim 10^{-3}$ cm-2 s-1 and $\zeta\simeq 10^{-16}\,$s-1, only a factor of a few below the observed rate shaded red). In addition, we also remark that there are other mechanisms that could boost the ionization rate without affecting the $511$ keV production: invoking more exotic DM scenarios, one can devise a scenario where asymmetric production of electrons and positrons (whilst maintaining charge conservation) can be achieved. We also remark that some known astrophysical sources could boost the ionization rate without affecting the $511$ keV production. Another intriguing possibility that may make both anomalies compatible with DM is that the direct interaction of DM particles with H2 molecules also produces ionization.
Figure 2: Left axis: Predicted $511$ keV bulge emission associated with the $e^{\pm}$ population from annihilations for various DM density profiles (solid lines). The shaded band shows the constraint set by the observed emission from the bulge, with conservative $50\%$ uncertainties. Right axis: Expected ionization rate in the CMZ (i.e. averaged within the CMZ) from annihilation of sub-GeV DM particles into $e^{+}e^{-}$, as obtained with the numerical calculation (dashed lines). The red band illustrates the range of observed values of ionization rate in the CMZ.
Since correlations between the $511$ keV line emission and the ionization rate in the CMZ can only happen if the source is injecting electrons or positrons simultaneously, we remark that observations of the ionization rate in the CMZ or even in molecular clouds, could help to discern the nature of the source producing the puzzling $511$ keV flux around the GC. However, it is clear that any source producing the emission of $511$ keV photons at the GC must also play a role in ionizing the CMZ. These correlations will be explored more thoroughly in a separate work. We also note that the broad range of empirical estimates of the molecular ionization rate (potentially favoring a relatively uniform distribution within the CMZ) makes it difficult to compare with the 511 keV line morphology, which suggests a cusped structure. Moreover, the angular resolution of the SPI detector (and similarly, the proposed COSI Siegert et al. (2020) experiment) remains too low (approximately $2.7^{\circ}$) to offer insights into the line’s distribution within the inner $\sim 200$ pc. Therefore, future observations with improved angular resolution are essential to better understand the emission within the CMZ.
Discussion and conclusions — In this Letter, we show that the anomalous ionization rate observed at the CMZ can be explained by DM particles of mass below a few tens of MeV annihilating into $e^{+}e^{-}$, for DM distributions with slopes slightly above that of the NFW $\gamma>1$. The thermally averaged annihilation cross-section required to match the observed ionization rates at the CMZ is below $10^{-33}$ cm3/s for a contracted NFW distribution ($\gamma=1.25$), and can be as low as $10^{-37}$ cm3/s for a Moore DM distribution, all well below the most stringent existing upper limits. Using both analytical approximate calculations and a more detailed numerical calculation with the DRAGON2 code, we show that the required annihilation cross-sections needed to explain the CMZ ionization rate are so low that annihilation of sub-GeV DM would produce no detectable inverse Compton, bremsstrahlung or synchrotron emission. Using extreme variations of our diffusion parameters, we find that uncertainties in the predicted ionization rate can be as high as an order of magnitude (see Section S.1.). However, given the low ionization rate required to explain the CMZ ionization rate and the fact that this mass region is largely unconstrained, these uncertainties can be easily compensated by a different value of the annihilation cross-section and provide the required ionization rate in the CMZ even for the most conservative of our estimations. While the power required to explain the ionization rate in the CMZ is around $10^{41}$ erg/s, DM following a NFW distribution would provide a power of $\sim 10^{44}\leavevmode\nobreak\ \rm erg/s$, in the case of a $1$ MeV particle and $\langle\sigma v\rangle\lesssim 10^{-33}$ cm3/s. In a Moore DM profile the power within the CMZ increases above $10^{50}\leavevmode\nobreak\ \rm erg/s$ for the same mass and annihilation cross-section. Furthermore, one of the strongest points of the DM explanation is that it produces a very flat radial profile of the ionization within the CMZ, as shown in the bottom panel of Fig. 1. For a NFW DM distribution, the radial profile of the ionization rate is quite flat, varying by no more than a factor of $3$ for a $5$ MeV DM particle, while much flatter at higher masses, for which the $e^{\pm}$ diffuse faster. For the Moore profile, representing the cuspiest DM distribution considered here, the ionization rate at different radial distances varies between a factor of a few and slightly more than an order of magnitude, values which are still compatible with the observations of the CMZ.
The fact that sub-GeV DM can even exceed the observed ionization rate in the CMZ without being in conflict with these probes demonstrates that this observable can be a very powerful tool for constraining sub-GeV DM. We find that the annihilation cross-section needed to explain the CMZ ionization rate is above, by at least a factor of a few, that needed to explain the bulge $511$ keV emission, whilst emphasizing that the uncertainties in the 511 keV emission make it difficult to presently rule out DM as the source of ionization in the CMZ. Therefore, we determine that the $511$ keV emission cannot conclusively rule out the DM explanation of the CMZ ionization rate for DM masses below a few MeV, leaving open a window to explore whether both observations are related (even beyond the DM hypothesis).
We conclude that the annihilation of sub-GeV DM that can explain the ionization rate in the CMZ could significantly contribute to an outstanding astrophysical mystery, namely the nature of the $511$ keV line emission. Finding spatial correlations between molecular clouds and positron line emission, as well as X-ray, $\gamma$ ray and radio constraints on correlated emissions may enable us to ascertain if the source of molecular ionization must be leptonic.
Acknowledgements— We would like to thank Marco Cirelli for helpful feedback on the draft of this letter. SB is supported by the STFC under grant ST/X000753/1. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the Swedish National Infrastructure for Computing (SNIC) under project No.2022/3-27 partially funded by the Swedish Research Council through grant no. 2018-05973.
Appendix A Propagation setup and calculations
As discussed in the main text, we used two different approaches to compute the steady-state diffuse distribution of electrons and positrons injected by DM in the CMZ. While the analytical approximation allows us easily estimate the average ionization within the CMZ, we need to use a numerical scheme to account for the effects of diffusion in the $e^{\pm}$. The details are given below.